# LLMs are not Zero-Shot Reasoners for Biomedical Information Extraction

### v1

@@ Introduction
Reasoning and knowledge enhancing techniques seem to not improve performance. Figure 1 and Figure  2 compare the results of the best performing techniques for each model for classification and NER, respectively. As seen in Table 1, perhaps counter-intuitively, Standard Prompting consistently achieves the highest average F1 scores across all models for classification task, with BioMistral-7B obtaining 36.48%, Llama-2-70B-Chat-AWQ achieving 40.34%, and Llama-2-7b-chat-hf scoring 34.92%. This result indicates that for structured prediction tasks, more complex reasoning techniques such as Chain of Thought (CoT) Prompting or Retrieval-Augmented Generation (RAG), do not outperform simpler approaches like Standard Prompting. For NER tasks, the results present a more nuanced picture compared to the classification tasks. While Standard Prompting remains effective, there is a noticeable shift in performance across different models and datasets. Notably, the scores are significantly lower than typical F1 scores in biomedical NER benchmarks. For instance, the NCBI disease corpus (Doğan, Leaman, and Lu 2014; Krallinger et al. 2015) and CHEMDNER dataset usually yield higher performances with specialized models or extensive pre-training. State-of-the-art models on these benchmarks can achieve Span F1 scores up to 0.90 for the NCBI disease corpus (Kocaman and Talby 2021; Zhou et al. 2023). However, similar to our findings, in true zero-shot setting, NER scores have been reported to be markedly low, even for the general domain (Shen et al. 2021) and when supplying label descriptions (Picco et al. 2024).
A possible reason for poor performance might be that these approaches have been tailored towards—and shown to work well on—knowledge- and reasoning-intensive tasks, such as Question Answering (Nori et al. 2023) or Mathematical Reasoning (Wang and Zhou 2024; Wang et al. 2022; Li et al. 2024). Meanwhile more narrowly defined tasks like information extraction or classification require the understanding of specific task semantics over generic reasoning capabilities. They seem to not require broad knowledge, as it could be found in biomedical paper abstracts or Wikipedia articles, but rather require application of domain knowledge in a specific and highly contextualized tasks, contained within the input document and task description. Models need to be able to handle highly specialized vocabulary, including jargon, acronyms, and synonyms that can vary widely between subfields (Kim et al. 2007; Zheng, Yu et al. 2018; Jiang and Xu 2024). There is a fundamental requirement for context dependent disambiguation of ambiguity and polysemy as well as nuances and variablity in syntax and expressions of biomedical concepts. This is often developed through specialized pre-training or domain-specific enhancements, which the LLMs have not been able to capture. These challenges necessitate models that not only have robust general NER capabilities but also an intricate understanding of biomedical context which can very for different subtasks within the domain.
@@
Scale drives improvements. In line with previous observations, we find that the 70B model also shows a considerable improvement (5.4% for classification, 2.2% for NER Span F1) over the 7B model. The most significant difference in performance between the Llama 7B and 70B Models is observed when using Self-Consistency with Chain of Thought and RAG (Wikipedia), where the 70B model outperforms the 7B model by 15.45% on classification and on NER tasks. This suggests that the larger model is significantly better at leveraging external knowledge when combined with self-consistency and chain of thought prompting. The larger model’s increased capacity might be particularly advantageous in handling these complexities, resulting in a more significant performance gap compared to simpler techniques. Methods like Chain-of-Thought Prompting and Self-Consistency with Chain-of-Thought and RAG involve complex reasoning and knowledge integration processes(Wei et al. 2022). This is further demonstrated by the fact that Llama 70B improves performance by 10.91% when using Self Consistency is added to Wikipedia based RAG, indicating that self consistency helps model combat the drop in performance when adding potentially irrelevant external information for the larger model. Unlike in classification tasks, where Standard Prompting was universally superior, NER performance does not degrade as much when using advanced prompting techniques, particularly when using larger models like Llama-2-70B, likely due to the general lack of epistemic confidence in the answers in the first place.
@@
Figure 5: Comparison of performance of each model in single label vs multi label datasets. Random baseline for single class classification is 0.415 and multi class classification is 0.215.
Detailed Comparison of Prompting Techniques
The use of CoT and Self Consistency are not helpful if there is a lack of parametric knowledge about the task. For BioMistral-7B, using Self-Consistency CoT prompting leads to the biggest reduction of about 16% for classification tasks. One possible reason is the domain-specific pre-training equips the model to better follow the instructions directly without needing additional reasoning structures, which seem detrimental. Similar to the RAG case, self-consistency seems to not consistently improve performance for NER. While Self Consistency aims to improve the reliability of Chain of Thought prompting by generating multiple reasoning paths and selecting the most consistent one, it might introduce additional complexity leading to errors or inconsistencies. This is especially true, if the model’s answers have low confidence scores due to insufficient parametric knowledge which prevents them to reliably solve these problems and would explain the observed performance drop. For NER tasks, the combination of Chain of Thought (CoT) and Self-Consistency prompting with RAG (Wikipedia) shows the most substantial performance difference between the 70B and 7B models. This suggests that larger models are more adept at leveraging external knowledge and complex reasoning strategies for entity recognition tasks if there is lack of parametric knowledge.
RAG does not help information extraction. The quality and relevance of the retrieved information can significantly impact performance, as seen from the fact that there is an average drop of 16.91% when using RAG with PubMed Corpora and 16.47% when using RAG with Wikipedia corpora as compared to the best performing technique for classification. While incorporating external knowledge through RAG can be generally beneficial for QA based tasks (Xiong et al. 2024) where incorporating relevant facts to the given question can append relevant knowledge into the model, it is not as straightforward in classification and information extraction tasks. This has to especially be considered in the given task setting, where the model could be confused by the presence of irrelevant knowledge information which adds an additional layer of complexity in extracting the relevant information for answering the relevant questions.
SC helps models filter out irrelevant noise in case of RAG, but does not help CoT While Self Consistency aims to improve the reliability of Chain of Thought prompting by generating multiple reasoning paths and selecting the most consistent one, is fundamentally dependent on the models epistemic certainty (Yadkori et al. 2024; Liu et al. 2024). This hinders performance if the model’s answers have low confidence scores due to insufficient parametric knowledge which prevents them to reliably solve these problems and would explain the observed performance drop. For BioMistral-7B, using Self-Consistency CoT prompting leads to the biggest reduction of about 16% for classification tasks. One possible reason is the domain-specific pre-training equips the model to better follow the instructions directly without needing additional reasoning structures, which seem detrimental. Similar to the RAG case, self-consistency seems to not consistently improve performance for NER. The combination of Chain of Thought (CoT) and Self-Consistency prompting with RAG (Wikipedia) shows the most substantial performance difference between the 70B and 7B models. This suggests that larger models are more adept at leveraging external knowledge and complex reasoning strategies for entity recognition tasks to augment the lack of epistemic uncertainty.
Detailed Per-dataset analysis
@@
Models Perform Significantly better on public datasets. Models perform significantly better on public datasets (average accuracy of 30%) compared to private datasets (average accuracy of 12%). This might hint at possible data leakage during pre-training or instruction-tuning, as publicly available datasets are more likely to be included in a web-crawl or a dedicated instruction tuning dataset. This might suggest that model performance on ‘unseen’ (yet publicly available) tasks could be a result of unintentional data leakage rather than a by product of reasoning or generalisation.
Multilingual Performance is not Scale Dependent. As shown in Figure 1, smaller models can match or even outperform larger models on Chinese and Japanese datasets but not on English datasets. This may be due to the heavy reliance on large English corpora during training, with limited exposure to medical contexts in other languages. This forces models to generalize compressed language representations to specialized domains, where overfitting on sparse languages may hinder larger models’ performance.
@@
LLMs struggle on tasks high complexity tasks As seen in Figure 5, LLMs seem to struggle to outperform random baselines for both single and multi class classification tasks. However, Figure 3 paints a more nuanced picture: guessing baseine remains unbeaten only on two of 14 datasets, which drags down the average performance significantly.
Figures 3 and 4 show that Llama2 70B demonstrates good performance in low-complexity tasks such as disease and symptom classification (CZIBase, NTCIR13-En) and medium-complexity tasks like Gene Expression classification (Geo). However, the model is challenged by higher-complexity problems, such as the BioNLP13-CG and GENIA-EE datasets.Specifically, in datasets that demand nuanced understanding and interpretation, such as the extraction of participants and outcomes from abstracts and gene ontology population (PICO, BioNLP13-GRO) the performance is low. When incorporating RAG (Retrieval-Augmented Generation) techniques, there are fluctuations in performance across datasets. While results improve on some datasets, RAG does not universally benefit the model’s ability to accurately extract and classify biomedical information.
Conclusion
@@
We provide a comprehensive benchmark and analysis of LLMs in Medical Classification and Named Entity Recognition tasks, revealing several key insights that have significant implications for the field. We carry out a critical investigation of broad claims regarding LLM capabilities by replicating them in various contexts, domains and datasets. We find that models suffer from fundamental drawbacks in generalizability, which hinder their performance in structured information extraction tasks on domain specific problems. This leads to Standard prompting outperforming more advanced methods across both the tasks. Our findings underscore the paramount importance of parametric knowledge capacity in zero-shot settings, regardless of advanced techniques used to augment external knowledge or model reasoning.
Acknowledgements
@@
This research is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. The authors thank Abhinav Ramesh Kashyap, Andy T. Liu and Vijay Prakash Dwivedi for their comments and useful feedback during the work. The authors further acknowledge and are thankful for the use of Imperial College Research Computing Service (DOI: http://doi.org/10.14469/hpc/2232), and the Computational Shared Facility at The University of Manchester.
Appendix A: Datasets
@@
Table 2 and 3 list the huggingface dataset cards and citations for each classification and ner dataset used in the paper respectively.
@@
Appendix A Appendix B: Compute Details
@@
GAD bigbio/gad (Bravo et al. 2015)
GEO bigbio/geokhoj_v1 (Elucidata 2022)
MedDialog bigbio/meddialog (Chen et al. 2020)
@@
LitCovid bigbio/bc7_litcovid (Chen et al. 2021)
CAS bigbio/cas (Grabar, Claveau, and Dalloux 2018)
ESSAI bigbio/essai (Grabar, Claveau, and Dalloux 2018)
NTCIR13-Ja bigbio/ntcir_13 _medweb (Iso et al. 2017)
NTCIR13-En bigbio/ntcir_13 _medweb (Iso et al. 2017)
NTCIR13-Zh bigbio/ntcir_13 _medweb (Iso et al. 2017)
PsyTAR bigbio/psytar (Zolnoori et al. 2019)
SciCite bigbio/scicite (Cohan et al. 2019)
@@
GeneTag-G bigbio/genetag (Tanabe et al. 2005)
GeneTag-C bigbio/genetag (Tanabe et al. 2005)
GENIA-PPI bigbio/genia _relation_corpus (Pyysalo et al. 2009; Hoehndorf et al. 2010; Ohta et al. 2010)
AnEm bigbio/an_em (Ohta et al. 2012)
BioInfer bigbio/bioinfer (Pyysalo et al. 2007)
Genia-$E$ bigbio/bionlp _shared_task_2009 (Kim et al. 2009)
BioNLP11-REL bigbio/bionlp_st _2011_rel (Pyysalo, Ohta, and Tsujii 2011)
BioNLP-13-CG bigbio/bionlp_st _2013_cg (Pyysalo, Ohta, and Ananiadou 2013)
BioNLP-13-GRO bigbio/bionlp_st _2013_gro (Kim et al. 2013)
BioNLP-13-PC bigbio/bionlp_st _2013_pc (Ohta et al. 2013)
PICO bigbio/ebm_pico (Nye et al. 2018)
MLEE bigbio/mlee (Pyysalo et al. 2012)

### v2

@@ 4.1 Overview of results
Reasoning and knowledge-enhancing techniques do not improve performance. Figures 1 and 2 compare the best-performing techniques for classification and NER. Surprisingly, Table 2 in the Appendix shows that Standard Prompting consistently achieves the highest average F1 scores across models: BioMistral-7B (36.48%), Llama-2-70B-Chat-AWQ (40.34%), and Llama-2-7b-chat-hf (34.92%). This suggests that for structured prediction tasks, complex reasoning techniques like CoT or RAG do not outperform Standard Prompting.
For NER, Standard Prompting remains effective, but performance varies across models and datasets. Scores are significantly lower than typical F1 scores in biomedical NER benchmarks such as NCBI disease corpus (Doğan et al., 2014; Krallinger et al., 2015) and CHEMDNER, where specialized models achieve up to 0.90 Span F1 scores (Kocaman and Talby, 2021; Zhou et al., 2023). However, similar to our findings, zero-shot NER scores tend to be low, even in general domains Shen et al. (2021) and when providing label descriptions Picco et al. (2024).
The likely reason for poor performance is that these approaches excel in knowledge- and reasoning-intensive tasks like Question Answering Nori et al. (2023) or Mathematical Reasoning Wang and Zhou (2024); Wang et al. (2022); Li et al. (2024), but structured prediction tasks require understanding task semantics over generic reasoning. These tasks rely less on broad knowledge from biomedical papers or Wikipedia and more on domain-specific application within the given input. Effective models must handle specialized vocabulary, jargon, acronyms, and synonyms varying across subfields Kim et al. (2007); Zheng et al. (2018); Jiang and Xu (2024). They must also resolve ambiguity, polysemy, and syntactic nuances in biomedical concepts, which the LLMs to not have been able to capture.
Scale drives improvements. Consistent with prior findings, the 70B model shows notable gains over the 7B model (5.4% for classification, 2.2% for NER Span F1). The largest performance gap appears when using SC with COT and RAG (Wikipedia), where the 70B model surpasses the 7B model by 15.45%. This suggests the larger model excels at leveraging external knowledge when paired with SC and chain-of-thought prompting. The 70B model’s greater capacity is particularly beneficial for handling complex reasoning and knowledge integration Wei et al. (2022). This is further supported by its 10.91% improvement when SC is added to Wikipedia-based RAG, helping mitigate performance drops from irrelevant external information. Unlike classification tasks, where Standard Prompting performed best, NER performance remains stable with advanced prompting techniques, especially in larger models like Llama-2-70B, likely due to the inherent lack of epistemic certainty in NER outputs.
@@
4.2 Detailed Comparison of Prompting Techniques
00.20.40.6L70BL7BBioF1 ScoreSingle-ClassMulti-Class
@@ 4.2 Detailed Comparison of Prompting Techniques
Figure 3: Performance comparison for BioMistral 7B, Llama7B and Llama70B on single- and multi-label datasets, with random guess baselines of 0.415 and 0.215, respectively.
CoT and SC underperform without sufficient parametric knowledge. For BioMistral-7B, SC-CoT prompts reduce classification performance by about 16%. One reason may be that domain-specific pre-training, while enhancing performance on specialized biomedical tasks, can limit the model’s general adaptability Brokman and Kavuluru (2024). Similar to RAG, SC does not consistently improve NER. While SC aims to generate multiple reasoning paths and select the most consistent one, it may introduce errors when the model’s epistemic certainty in its outputs is low, leading to performance drops. For NER, combining CoT and SC with RAG (Wikipedia) produces the largest performance gap between 70B and 7B models, suggesting that larger models use external knowledge and complex reasoning more effectively when parametric knowledge is limited.
RAG does not help information extraction. Although RAG can improve QA tasks by retrieving relevant facts Xiong et al. (2024), it seems less useful for classification and information extraction, where irrelevant information misleads the model, creating additional complexity. This results in a drop in classification accuracy, dropping 16.91% with PubMed Corpora and 16.47% with Wikipedia compared to the best classification method.
SC helps filter noise for RAG but does not aid CoT. While SC aims to improve CoT by generating multiple reasoning paths, its efficacy depends on the model’s epistemic certainty Yadkori et al. (2024); Liu et al. (2024). Insufficient parametric knowledge lowers confidence, causing performance declines. BioMistral-7B loses about 16% in classification with SC-CoT prompting. SC also fails to consistently boost NER. However, in the 70B model, combining CoT and SC with RAG (Wikipedia) yields notable gains, indicating that larger models better exploit external knowledge and present higher epistemic certainty owing to their larger parametric capacity.
4.3 Detailed Per-Dataset Analysis
Figure 3 provides the performance comparison of the three models on single and multi-label tasks. Figures 4 and 5 provide a detailed analysis and breakdown of performance of each technique (NER and classification) on each dataset, along with random baselines for each. We discuss their implications below.
@@ 4.3 Detailed Per-Dataset Analysis
Models perform significantly better on public datasets. On public datasets, models average around 30% accuracy, compared to 12% on private datasets, suggesting potential data leakage from publicly available sources used in pre-training or instruction-tuning. Thus, performance on “unseen” tasks may stem from prior exposure rather than true generalization.
Multilingual performance is not scale dependent. As shown in Figure 1, smaller models can match or outperform larger ones on Chinese and Japanese datasets but lag behind in English tasks. This disparity likely results from heavy reliance on English corpora and limited non-English domain exposure, increasing overfitting risks. Factors like language family, data sampling ratios, and sparse representation can also inhibit multilingual models He et al. (2024); Bagheri Nezhad and Agrawal (2024).
@@
LLMs struggle on high-complexity tasks. As Figure 3 shows, on average, the models fail to surpass random baselines for single and multi-class classification, though Figure 4 shows these baselines remain unbeaten in only two of 14 datasets.
5 Conclusion
@@
We provide a comprehensive benchmark and analysis of LLMs in Medical Classification and Named Entity Recognition tasks, revealing several key insights that have significant implications for the field. We carry out a critical investigation of broad claims regarding LLM capabilities by replicating them in various contexts, domains and datasets. We find that models suffer from fundamental drawbacks in generalizability, which hinder their performance in structured information extraction tasks on domain specific problems. This leads to standard prompting outperforming more advanced methods for all models. Our findings underscore that parametric knowledge capacity remains the most important factor in zero-shot settings, with advanced techniques used to augment external knowledge or model reasoning failing to improve performance.
Acknowledgement
@@
Viktor Schlegel is part of the IN-CYPHER programme and is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. We are grateful for the support provided by Research IT to him in form of access to the Computational Shared Facility at The University of Manchester and the computational facilities at the Imperial College Research Computing Service [^1] DOI: https://doi.org/10.14469/hpc/2232.
Limitations
While our study provides important insights into LLMs’ capabilities for biomedical classification and information extraction, several limitations should be considered when interpreting our results. Our findings are primarily empirical and, although they suggest consistent patterns across models and tasks, further theoretical work is needed to fully understand why advanced prompting techniques fail to improve performance on structured prediction tasks. We deliberately exclude closed-source LLMs from our analysis due to privacy concerns in medical applications and the observed dataset leakage issues, where public dataset contamination is even harder to control for proprietary models. Additionally, our focus on constrained decoding for reliable output parsing necessarily limits us to open-source models where we have access to the generation process.
We also specifically choose not to evaluate in-context learning (ICL) approaches, as our study focuses on “true” zero-shot capabilities where no task-specific examples are available. While techniques like k-NN ICL have shown promise in other domains, they require substantial annotated data to retrieve examples from—which is often unavailable in practical medical settings. Fixed ICL examples could be used, but performance would then largely depend on example selection, essentially reducing the evaluation to the quality of prompt engineering. To balance (i) scientific validity and focus on real-world scenarios, where domain experts may not be prompt engineering specialists, with (ii) the need to provide useful information to the models, we instead opt for the zero-shot setting—addressing (i)—while ensuring semantic clarity through meaningful label names (e.g., using “Control” and “Perturbation” rather than “0” and “1” in the GEO dataset)—addressing (ii).
Appendix A Related Work
Existing LLMs Benchmarks: With the rising popularity of LLMs, many works evaluated their performance in the biomedical and clinical domains. These works typically focus on evaluating domain-knowledge by means of Question Answering Singhal et al. (2023); Harris (2023); Subramanian et al. (2024), or focus directly on possible application scenarios, such as summarisation Li et al. (2023); Yim et al. (2023) or clinical coding Kaur et al. (2023). Many works combine these two directions in an effort to provide more comprehensive benchmarks Srivastava et al. (2024); Xiong et al. (2024); Feng et al. (2024); Chen et al. (2020); Manes et al. (2024). However, many of these works overlook the wealth of existing literature and plethora of available resources for traditional structured prediction tasks in the biomedical domain, such as document classification, entity recognition and linking and event and relation extraction (e.g., Pyysalo et al. (2007; 2012) to name a few). Fries et al. (2022a) have provided a comprehensive and unified collection of these resources, however their work prioritises reportage of the resource collection over benchmarking results. Their preliminary evaluations suggest that their evaluated pre-LLM era models barely surpass the random guess baseline in the zero-shot setting. We build upon their work by providing a detailed analysis to what extent approaches to enhance reasoning and knowledge in LLMs help to challenge this status quo.
Appendix B Datasets
@@
Table 3 and 4 list the huggingface dataset cards and citations for each classification and ner dataset used in the paper respectively.
@@
Appendix C Compute Details
@@
GAD bigbio/gad Bravo et al. (2015)
GEO bigbio/geokhoj_v1 Elucidata (2022)
MedDialog bigbio/meddialog Chen et al. (2020)
@@
LitCovid bigbio/bc7_litcovid Chen et al. (2021)
CAS bigbio/cas Grabar et al. (2018)
ESSAI bigbio/essai Grabar et al. (2018)
NTCIR13-Ja bigbio/ntcir_13 _medweb Iso et al. (2017)
NTCIR13-En bigbio/ntcir_13 _medweb Iso et al. (2017)
NTCIR13-Zh bigbio/ntcir_13 _medweb Iso et al. (2017)
PsyTAR bigbio/psytar Zolnoori et al. (2019)
SciCite bigbio/scicite Cohan et al. (2019)
@@
GeneTag-G bigbio/genetag Tanabe et al. (2005)
GeneTag-C bigbio/genetag Tanabe et al. (2005)
GENIA-PPI bigbio/genia _relation_corpus Pyysalo et al. (2009); Hoehndorf et al. (2010); Ohta et al. (2010)
AnEm bigbio/an_em Ohta et al. (2012)
BioInfer bigbio/bioinfer Pyysalo et al. (2007)
Genia-$E$ bigbio/bionlp _shared_task_2009 Kim et al. (2009)
BioNLP11-REL bigbio/bionlp_st _2011_rel Pyysalo et al. (2011)
BioNLP-13-CG bigbio/bionlp_st _2013_cg Pyysalo et al. (2013)
BioNLP-13-GRO bigbio/bionlp_st _2013_gro Kim et al. (2013)
BioNLP-13-PC bigbio/bionlp_st _2013_pc Ohta et al. (2013)
PICO bigbio/ebm_pico Nye et al. (2018)
MLEE bigbio/mlee Pyysalo et al. (2012)
@@
Appendix D Methodology
Datasets
Since we evaluate different prompting techniques, we restrict the choice of tasks to those where the number of possible labels is small enough to fit in the evaluated LLMs’ context window. We restrict the number of labels to ten and the mean length of the input documents to at most 2048 tokens. This leaves us with 14 different classification datasets from the BigBio collection222for the GAD dataset, we only select 1 fold out of the 10 available, as the folds feature the same task for different data, unlike other datasets. We also skipped the Chinese subset of meddialog as we had difficulties loading the dataset. For the NER task, we sample 12 datasets from the pool of those that satisfy the criteria. The resulting dataset sample features four non-English datasets and six non-public classification datasets, which allows us to investigate whether LLMs perform better on minority languages or on data that is less likely to be found in public pre-training corpora. We run the evaluation on the official test-set split where available, otherwise we consider the full dataset. For datasets with more than 500 instances, we sample 500 random but fixed instances to speed up the experiments. Overall, our selection spans English and non-english source data, publicly available and private datasets, and various domains such as scientific papers, medical notes and social media.
Classification: The datasets used for classification tasks include both single-label and multi-label datasets, covering a wide range of biomedical and clinical domains. For single-label classification, the GAD dataset focuses on identifying associations between genes and diseases Bravo et al. (2015), while the GEO dataset is concerned with classifying microarray, transcriptomics, and single-cell experiments from the Gene Expression Omnibus (GEO) database Elucidata (2022). The MedDialog dataset aims to classify dialogue snippets as either being said by a doctor or a patient Chen et al. (2020). Furthermore, the CZIDrsm dataset has several subsets, including one for classifying research articles based on aspects of disease research (CZIBase), and others for identifying whether a paper describes substantive research into Quality of Life (CZIQoL) or is a natural history study (CZINatHist).
In multi-label classification, the LitCovid dataset is used for the classification of COVID-19-related articles Chen et al. (2021). The CAS and ESSAI datasets are utilized for identify negation and uncertainty clinical cases from French-speaking countries Grabar et al. (2018). The NTCIR13 datasets include subsets for disease classification of tweets in Japanese (*-Ja), English (*-En), and Chinese (*-Zh) Iso et al. (2017). Additionally, the PsyTAR dataset is used for sentence classification of various drug-related effects, such as Adverse Drug Reactions (ADR) and Withdrawal Symptoms (WDs) Zolnoori et al. (2019), while the SciCite dataset is used for citation intent classification based on the context within computer science and biomedical domains Cohan et al. (2019).
NER: The datasets for Named Entity Recognition (NER) tasks are similarly divided into entity recognition (single entity type) and classification (multiple entity types). In the single-type category, the GeneTag dataset is used for gene/protein NER, with two annotation versions: the original GeneTag-G and the corrected GeneTag-C Tanabe et al. (2005). Additionally, the GENIA-PPI dataset focuses on protein-protein interactions or gene regulatory relations within the GENIA corpus, capturing primarily static relations Pyysalo et al. (2009); Hoehndorf et al. (2010); Ohta et al. (2010).
The multiple-type NER datasets encompass various complex biomedical tasks. The AnEm dataset targets anatomical entity recognition Ohta et al. (2012), while the BioInfer dataset focuses on recognizing proteins, genes, and RNA entities Pyysalo et al. (2007). The Genia-EE dataset is used for the GENIA Event corpus Kim et al. (2009), and the BioNLP11-REL dataset is employed for extracting part-of relations between genes/proteins and associated entities Pyysalo et al. (2011). Furthermore, the BioNLP-13-CG dataset is used for Cancer Genetics (CG) information extraction, focusing on recognizing events represented as structured n-ary associations of given physical entities Pyysalo et al. (2013). The BioNLP-13-GRO dataset aims to populate the Gene Regulation Ontology with events and relations Kim et al. (2013), and the BioNLP-13-PC dataset is used for the automatic extraction of biomolecular reactions from text Ohta et al. (2013). Lastly, the PICO dataset deals with recognizing (P)articipants, (I)nterventions, and (O)utcomes Nye et al. (2018), and the MLEE dataset is used for event extraction related to angiogenesis Pyysalo et al. (2012).
Models
For our experiments, we employed two instruction-tuned variants of the Llama-2 model—7B and 70B—both Touvron et al. (2023), alongside the BioMistral-7B model Labrak et al. (2024) which was further pre-trained on the biomedical domain. Since we make use of constrained generation to generate model outputs and guide the models decoding process, we retrict the evaluation to open source models since this process is not possible for proprietary models such as GPT-4.
Techniques
Table 1 summarizes the techniques used in this study and highlights relevant nuances and comments. These techniques include Vanilla (standard prompting), CoT (chain-of-thought reasoning) Wei et al. (2022), and SC (self-consistency) Wang et al. (2022), as well as RAG (retrieval-augmented generation) Lewis et al. (2020). For RAG, we used FAISS Douze et al. (2024); Johnson et al. (2019) with PubMed abstracts Sanyal et al. (2021) and Wikipedia articles as corpora, embedding documents with all-MiniLM-L6-v2 Reimers and Gurevych (2019). We also implemented constrained decoding for structured output generation Willard and Louf (2023), crucial for ensuring reliable outputs in NER and classification tasks. A novel two-stage approach for NER was adopted, inspired by Shen et al. (2021), where general entities were assigned in Stage 1 and refined in Stage 2.
Standard prompting was used as a baseline for both the Classification as well as the NER tasks. Chain-of-thought reasoning Wei et al. (2022) has been shown to improve performance, particularly in QA and logical reasoning tasks. Thus, we also ran experiments with chain-of-thought reasoning to measure its impact on model performance. For the NER task, we adapted a more guided, two-stage approach Shen et al. (2021) to implement a novel chain-of-thought reasoning approach. Here, The first stage involves inducing a generic entity name from a datasets’ known entity labels—e.g., “Bodypart” for the NER labels describing different bodyparts—and then labelling the input document with that generic entity type. In the second stage all entities labelled in this way are further disambiguated with their respective fine-grained dataset NER labels. Retrieval Augmented Generation Lewis et al. (2020) has been established as an effective technique to improve model performance by introducing relevant non-parameteric knowledge to models and thus grounding the generated outputs to factual information. Xiong et al. (2024) conducted a systematic study of RAG on medical QA, and we incorporate their findings into our study. We used PubMed abstracts Sanyal et al. (2021) and Wikipedia articles as knowledge corpora, because Xiong et al.’s (2024) experiments found that using PubMed improved performance over non RAG techniques, while using Wikipedia reduced performance in medical QA tasks. Our goal was to evaluate whether the same holds true for structured prediction tasks as well. For the RAG module, we made use of FAISS Douze et al. (2024); Johnson et al. (2019), which allows retrieval of most similar documents based on semantic similarity, where we used the all-MiniLM-L6-v2 sentence transformers Reimers and Gurevych (2019) model for embedding input documents and corpora. For each experiment, the number of retrieved documents was computed based on the maximum possible documents which could be used without exceeding the token limit of the model.
Self-consistency, proposed by Wang et al. (2022), improves chain-of-thought reasoning of LLMs by sampling reasoning paths for a given problem, followed by a majority vote for the final answer. We also conduct a set of experiments employing self-consistency to investigate whether such improvements can be observed on structured prediction tasks in the medical domain as well. For classification tasks, self consistency was employed to generate multiple reasoning chains for the given problem, followed by answer extraction from each reasoning chain and majority voting to select the final answer. For NER tasks, since we follow the two-stage approach, self-consistency was employed in both stages. Multiple general entity labels were generated in the first stage, and entities were extracted for each such label. In the second stage, self consistency was again used for the entity selection phase as well as the entity label determination step. Majority voting was utilised in final label or class selection in each case Xie et al. (2023).
Constrained decoding in LLMs Willard and Louf (2023) was used to ensure structured information extraction and text generation. This allowed us to evaluate the LLMs for the task at hand without the added variability due to the aleatoric uncertainties brought about by the probabilistic language generation fundamental to the architectures of the models. More specifically, for classification tasks, we ensured the presense of at least one label in the generated outputs. For NER we restricted the generation of spans occurring in text in the first step, and in the second step, for each of the spans we restricted the generation to any of the possible labels. This is also one of the reasons why we opted against evaluating API-based closed-source LLMs [^3] The other reason being their intransparancy with regard to training data, which violates our “true” zero-shot setting., as in our initial experiments the hallucinations in generated outputs created problems with reliably parsing the structured outputs.
Appendix E Analysis and Performance Breakdown
Figures 4 and 5 provide a detailed analysis and breakdown of performance of each technique (NER and classification) on each dataset, along with random baselines for each. Figure 3 provides the performance comparison of the three models on single and multi-label tasks. A complete discussion for these figures and their implications can be found in section 4.3.
As discussed in section 4.3, LLMs struggle on high-complexity tasks. Even the best performing model, Llama2 70B performs well on only relatively low-complexity tasks (CZIBase, NTCIR13-En) and moderate tasks (Geo), but struggles with higher-complexity datasets (BioNLP13-CG, GENIA-EE). In tasks requiring nuanced interpretation (PICO, BioNLP13-GRO), performance remains low. Although RAG (Retrieval-Augmented Generation) sometimes boosts results, it does not universally enhance biomedical information extraction or classification. These findings indicate that even the most advanced general-purpose and domain-specific LLMs are not good zero-shot reasoners for structured prediction tasks such as biomedical information extraction, especially for complex task settings.
Appendix F Results Analysis